OpenStudy (anonymous):

Propagated error? The measurements of the base and altitude of a triangle are found to be 36 and 50 centimeters, respectively. The possible error in each measurement is 0.25 centimeter. Use differentials to approximate the possible propagated error in computing the area of the triangle.

OpenStudy (anonymous):

I have to find the derivative of the volume equation or something and set delta V = dv... right?

OpenStudy (anonymous):

What throws me off is that I have to account for two variables having calculated error

OpenStudy (anonymous):

dA = dA/da da + dA/db db a = 50 b = 36 dA/da = b/2 = 36/2 = 18 dA/db = a/2 = 50/2 = 25 da = db = 0.25 So the possible propagated error is given by dA = 18 * 0.25 + 25 * 0.25 = 4.5 + 6.25 = 10.75 cm^2 This stuff makes my head hurt! Dont ever ask me to do c alclus again! This goes for all of you!! :) lol jk Have fun understanid it!

OpenStudy (anonymous):

I saw that on Yahoo Answers... like, why is dA/da = b/2 and dA/db = a/2?

OpenStudy (anonymous):

Because your calculating basically the same thing every time, because your using the same margin for error. http://www.youtube.com/watch?v=R7t9Iv_lTc8 watch this tutorial

OpenStudy (anonymous):

Thanks :) can you tell I hate math passionately lol